5. Real wages, nominal wages, and unexpected changes in the price level


Deborah currently earns a nominal - wage of $ 12.00 per hour; in other words, the amount of her paycheck each week is $ 12.00 per hour times the number of hours she works. Suppose the price of sparkling water is $ 2.50 per gallon; in this case, Deborah's real [ wage, in terms of the amount of sparkling water she can buy with her paycheck, is 6 gallons of sparkling water per hour.


When workers and firms negotiate compensation packages, they have expectations about the price level (and changes in the price level) and agree on a nominal - wage with those expectations in mind. If the price level turns out to be higher than expected, a worker's real ∇ wage is lower ∇ than both the worker and employer expected when they agreed to the wage.


Deborah and her employer both expected inflation to be 4 % between 2012 and 2013, so they agreed, in a two-year contract, that she would earn $ 12.00 per hour in 2012 and $ 12.48 per hour in 2013 . However, suppose inflation between 2012 and 2013 actually turned out to be 7 %, not 4 %. For example, suppose the price of sparkling water rose from $ 2.50 per gallon to $ 2.68 per gallon. This means that between 2012 and 2013, Deborah's nominal wage by %, and her real wage by approximately
Answer #1

Blanks:-

1- Nominal

2-Real

3 -12/2.50 = 4.8

4- Nominal

5- Real

6- Lower

7- Nominal wage increases by 12.48-12/12*100 = 4%

8) Real wage decreases by(12.48/2.68) -4.8/4.8*100)) = -3%